# Dynamic borrowing from historical controls via the synthetic prior with covariates in randomized clinical trials

### v1

@@ 6 Discussion
More generally, trialists wanting to prospectively calibrate the borrowing/no-borrowing tradeoff (i.e. power/Type I error tradeoff) for a specific planned trial may find the best success in tuning one of two method parameters. First, they may experiment with increasing or decreasing $p_{ind}$, the prior probability of the no-borrowing submodel, as we do in Table S1. Alternatively, they may experiment with changing the decision rule threshold $q_{positive}$ or $q_{clinical}$ for the trial’s final treatment effect inference in equations (2) or (3), which would not require new simulation for each value if MCMC output has been saved. Both reducing $p_{ind}$ and $q_{clinical}$ or $q_{positive}$ would reduce Type I error at the cost of also reducing power.
@@
Because the method only requires group-level data, it may be possible to include trials where the patient-level data would not be available due to privacy or intellectual property concerns. The group-level covariates might ideally be believed to be strong predictors based on solid theoretical or past empirical evidence, though in settings with enough historical trials it may be possible to incorporate higher-dimensional covariates through the use of sparsity-inducing priors on the regression coefficients. While not the focus of the present work, using the SPx modeling strategy with patient-level data and covariates would likely reduce trial sizes and increase accuracy even more substantially. The proposed SPx model could easily be modified to accommodate the patient-level data. For example, notation $y_{hj}$ and $x_{hj}$ would represent the response and covariates for patient $j$ from study $h$. Then patient-specific random effects $\psi_{hj}$ may be introduced and a prior model $f(\psi_{hj}|\psi_{h})$ may be used to allow shrinkage of the random effects to study-specific effect $\psi_{h}$. The SPx model can then be used for $\psi_{h}$ to complete the model construction.

### v2

@@ 6 Discussion
Like the RMAP method, SPx uses Bayesian model averaging to adaptively balance between borrowing from the historical data and discounting it when appropriate. However, we highlight two important differences between SPx and RMAP. First, because SPx includes both a direct-borrowing component and a regression component, it will be able to robustly gain accuracy whenever either (a) the new trial’s control rate is similar to the historical control rates or (b) group-level covariates predict the response rates. In contrast, RMAP only gains accuracy in case (a). And second, SPx can be more robust than RMAP to different types of discrepancies between the historical and new control data, especially for smaller trials. See for example Table 2, where SPx has better control of Type I error rate than RMAP in Scenarios 2 and 4 when $n^{c}_{max}=100$. This is due to the careful construction of the SPx hyperprior (i.e. its borrowing submodels have concentrated, aggressive priors but the prior model weights favor the independent submodel), as we discuss in Sections 2.2 and 4.3.
@@
More generally, trialists wanting to prospectively calibrate the borrowing/no-borrowing tradeoff (i.e. power/Type I error tradeoff) for a specific planned trial may find the best success in tuning one of several key parameters of the method. First, they may experiment with increasing or decreasing $p_{ind}$, the prior probability of the no-borrowing submodel, as we do in Table S1. They may also consider making $p_{hist}$ and $p_{reg}$ unequal if one of these submodels is less suited for scenarios they are concerned about. Alternatively, they may experiment with changing the decision rule threshold $q_{positive}$ or $q_{clinical}$ for the trial’s final treatment effect inference in equations (2) or (3), which would not require new simulation for each value if MCMC output has been saved. Both reducing $p_{ind}$ and $q_{clinical}$ or $q_{positive}$ would reduce Type I error at the cost of also reducing power. Further, it may be desirable to tune the values $n^{c}_{max}$, $\gamma_{max}$, and $\gamma_{min}$ in the adaptive design in order to achieve specific practical constraints or operating characteristics.
@@
Because the method only requires group-level data, it may be possible to include trials where the patient-level data would not be available due to privacy or intellectual property concerns. The group-level covariates might ideally be believed to be strong predictors based on solid theoretical or past empirical evidence, though in settings with enough historical trials it may be possible to incorporate higher-dimensional covariates through the use of sparsity-inducing priors on the regression coefficients. We note that because we use trial-level summary statistics as covariates, one might be concerned that variability (i.e. measurement error) in these statistics could affect the performance of SPx. However, SPx uses these covariates only to improve prediction of the new trial’s control response rate via the linear predictor $\bm{\beta}^{\top}\bm{x}_{H+1}$. It is well-established that while measurement error in covariates can bias estimation of regression coefficients [Carroll 2006], it does not generally reduce prediction accuracy (Carroll [2006], Chapter 2.6; Khudyakov et al. [2015]). For this reason, we do not expect measurement error in the group-level covariates to substantially impair SPx’s predictions. An important exception arises if the magnitude or structure of measurement error differs substantially between the historical and new trials (Carroll [2006], Chapter 2.6; Luijken et al. [2019]), and this is an important area for future work. In this case, it would be necessary to explicitly model the measurement errors.
While not the focus of the present work, using the SPx modeling strategy with patient-level data and covariates would likely reduce trial sizes and increase accuracy even more substantially. The proposed SPx model could easily be modified to accommodate the patient-level data. For example, notation $y_{hj}$ and $x_{hj}$ would represent the response and covariates for patient $j$ from study $h$. Then patient-specific random effects $\psi_{hj}$ may be introduced and a prior model $f(\psi_{hj}|\psi_{h})$ may be used to allow shrinkage of the random effects to study-specific effect $\psi_{h}$. The SPx model can then be used for $\psi_{h}$ to complete the model construction. We also note that with patient-level data it may be feasible for analysts to use more sophisticated approaches to control for covariates (compared to our linear regression of trial-level covariates over a moderate number $H+1$ of studies). For example, with patient-level data one might consider replacing SPx’s regression submodel with a semiparametric model that allows flexible covariate-response relationships [Zhou and Ji 2021], a model that integrates propensity scores [Lin et al. 2018, Chen et al. 2022], or other semiparametric Bayesian strategies [Müller et al. 2011].